OpenStudy (anonymous):

At Sanger’s auto garage, 40% of the cars brought in for service need an oil change. Of the cars that need oil changes, 30% also need a tire rotation. Find the probability that a car that comes into the auto garage needs both an oil change and a tire rotation.

OpenStudy (anonymous):

A. 0.10 B. 0.12 C. 0.70 D. 0.75

OpenStudy (bibby):

In probability of usually means multiplication. What is 30% of 40%?

OpenStudy (anonymous):

Im doing conditional probability

OpenStudy (bibby):

That isn't the right answer chaz

OpenStudy (anonymous):

0.75

OpenStudy (bibby):

how'd you get that?

OpenStudy (anonymous):

?

OpenStudy (bibby):

I'm not going to give you the answer.

OpenStudy (bibby):

Provide your work and logic

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

im not understanding conditional geometry...

OpenStudy (anonymous):

P(Oil Change)=.4 P(Tire Rotation|Oil Change)=.3 P(A|B) = P(A and B)/P(B)

OpenStudy (anonymous):

.75 is the correct answer

OpenStudy (anonymous):

but its not :(

OpenStudy (anonymous):

im taking a quiz and it said it was incorrect :(

OpenStudy (anonymous):

oops, i mean .12

OpenStudy (anonymous):

ok thanks :D

OpenStudy (bibby):

30% of the 40% that require the thing = 0.3*0.4

OpenStudy (bibby):

Thanks for the explanation. You people are ruining this site

OpenStudy (anonymous):

.3=P(A and B)/.4 => P(A and B) = .3*.4 = .12

OpenStudy (anonymous):

does that make sense?